A boy stands on a diving board and throws a stone into a swimming pool. The stone is thrown from a height of 3.67 m above the water surface with a velocity of 5.35 m/s at an angle of 60.0° above the horizontal. Determine the horizontal distance the stone travels before it strikes the water. Answer: Check
Answer #1
vxo = vo·cos θ 

vyo = vo·sin θ
vy = vyo - g·t

trise = vyo/g
h = yo + vyo·t - 0.5·g·t2
0 = h - 0.5·g·t2

tfall = √(2·h/g)

The total flight time of the projectile is then:

tflight = trise + tfall
range = vxo·tflight

range = 2.675 x 1.3373 = 3.577 m

please rate it up thanks :)